# Pharos-ESG: A Framework for Multimodal Parsing, Contextual Narration, and Hierarchical Labeling of ESG Reports

### v1

@@ Introduction
These challenges have driven financial research to adopt indirect proxies for ESG content, such as simple disclosure indicators (Ni2019Mandatory), small-scale case studies (Berg2022AggregateConfusion), or third-party ratings (Christensen2022doclaynet), all of which bypass the rich semantics of the reports themselves. Meanwhile, existing document parsers, designed for structurally regular formats like academic papers, legal contracts, or forms (Palm2019Attend; Xu2020layoutlm; Huang2022LayoutLMv3), struggle with the irregularity and implicit structure of ESG reports, limiting their applicability to this domain.
To address these challenges, we present Pharos-ESG, a unified framework for multimodal parsing, contextual grounding, and hierarchical labeling of ESG reports (Figure 2). Pharos-ESG integrates four core components: (a) reading-order modeling based on layout flow, (b) structure reconstruction guided by table-of-contents (ToC) anchors, (c) contextual transformation of visual elements into natural language, and (d) multi-level labeling across ESG topics, GRI indicators, and sentiment. The system outputs reading-order-aligned structures with linked text, tables, and images, enriched by hierarchical and semantic annotations. Experiments on expert-annotated benchmarks demonstrate that Pharos-ESG consistently outperforms both specialized parsers (e.g., MinerU (Huang2022LayoutLMv3), Marker (Tito2023Hierarchical), Docling (Kim2023Donut)) and general-purpose multimodal models (e.g., GPT-4o, Gemini 2.5 Pro, Doubao) across multiple tasks.
@@
Recent advances in document intelligence have significantly improved performance on general-purpose document understanding tasks (Xu2020layoutlm; Huang2022LayoutLMv3). However, ESG reports remain challenging due to chaotic reading order from irregular, slide-like layouts and implicit hierarchies caused by lengthy, weakly structured content.
@@
Dedicated document parsing systems such as Docling (livathinos2025docling), MinerU (wang2024mineru), Marker (marker2025), and TextIn (textin2023), are effective at converting documents into structured outputs. They often incorporate state-of-the-art layout analysis models like EfficientViT (liu2023efficientvit), RT-DETR (zhao2024detrs), and LayoutLMv3 (Huang2022LayoutLMv3), achieving high accuracy on datasets with regular layouts such as PubLayNet (zhong2019publaynet). However, their performance degrades significantly on structurally diverse datasets like DocLayNet (pfitzmann2022doclaynet), revealing limitations in generalization to visually irregular documents, including ESG reports. In parallel, vision-language models (VLMs) such as GPT (OpenAI), Gemini (Google), and DeepSeek (DeepSeek) provide a flexible, end-to-end solution by directly converting page images into text. However, these models often hallucinate and incur substantial computational costs when reconstructing implicit hierarchies in long, weakly structured ESG reports.
@@
Transformer-based models such as DocFormer (Appalaraju2021DocFormer), LayoutLMv3 (Huang2022LayoutLMv3), StrucTexT (Appalaraju2021DocFormer), UDoc (Kim2023Donut), and SelfDoc (Li2021SelfDoc) employ cross-modal attention to associate figures with captions and link textual blocks with related tables or legends. While effective on structurally regular documents, they struggle with ESG reports, where semantically related elements are often dispersed across non-adjacent, even cross-page regions.
@@
Traditional approaches typically model structure at the page level, limiting global context and failing to capture cross-page semantic and hierarchical links (Huang2022LayoutLMv3; Xia2022Structured). Most rely on explicit visual cues (e.g., numbering, indentation, font size) which often prove unreliable in complex or inconsistent layouts (Wang2023DocParser; zhong2019publaynet; Xu2020layoutlm). Recent models combine multimodal encoders with structure-aware decoders for joint hierarchy prediction, such as HRDoc (Marulanda2023Correspondence) and DINO (Zhang2023DINO). Yet, attention decay in lengthy documents (e.g., 20+ pages) still leads to recognition failures (Beltagy2020Longformer; Huang2022LayoutLMv3). These challenges are amplified in ESG reports, which feature exceptional length, weak structural cues, and diverse formatting across issuers.
@@
To our knowledge, four recent datasets support ESG-related research. morio2023nlp released 10,000+ corporate climate policy documents. ESG-FTSE (wang2024esg) contains 3,913 news articles on the UK’s top 10 firms. A3CG (ong2025towards) compiles ESG reports from 1,679 Singapore-listed companies, and DynamicESG (tseng2023dynamicesg) includes 2,220 Taiwan-based ESG news articles. However, these datasets offer limited coverage of full-length, multimodal ESG reports and lack the fine-grained annotations required for downstream financial tasks.
@@
Based on the extraction of multimodal elements from ESG reports via a structured preprocessing pipeline—including metadata extraction (han2003automatic), layout analysis (Huang2022LayoutLMv3), content parsing (zhang2020trie), and noise reduction (easyocr2024)—Pharos-ESG adopts and extends the Immediate Succession During Reading paradigm (zhang2024modeling) for reading order modeling, transitioning from global sequence ranking to a pairwise succession classification framework.
@@
where $E(w_{i})$ and $E(w_{j})$ are content embeddings: text and tables are encoded using LayoutLMv3 (Huang2022LayoutLMv3) For image blocks, the local image URL is used only as a placeholder for block identity and is not encoded nor used as a visual feature in the model. $\Delta y_{ij}$ and $\Delta x_{ij}$ denote center offsets; $\mathrm{IoU}{ij}$ and $\mathrm{Dist}{ij}$ are the bounding-box overlap and normalized distance; and $E(c_{i}),E(c_{j})$ are one-hot encodings of block types.
@@
To align the hierarchical structure parsed from ToC with the corresponding content in the document body, we propose ALIGN (Anchor-based Linguistic Indexing for Granular Navigation), a multi-stage alignment algorithm designed for visually heterogeneous layouts. ALIGN operates in three stages: (1) exact match, performing character-level matching to identify high-confidence anchors; (2) fuzzy/containment matching, using Levenshtein similarity (Mehlhorn1990Dynamic) and substring containment to expand coverage; and (3) context-aware insertion, resolving unmatched ToC headings by reasoning within anchor-defined windows in the document body. To this end, a large language model is prompted using a tailored CIP (Context-aware Insertion Prompt) to perform four structured inference steps: (i) summarizing paragraphs to generate semantic representations; (ii) assessing whether the region lacks an overview heading; (iii) identifying where the unmatched heading improves local structure; and (iv) selecting the optimal insertion point based on structural and semantic alignment.
@@
where $h_{i}$ is the hierarchical heading path, $\{x_{1},\ldots,x_{k}\}$ are the ordered elements, and $q_{i}$ is the task instruction. The prompt comprises (i) heading path embedding, linking the cluster to its section context, and (ii) element declaration templates, specifying type, order, and content. In the multimodal embedding and semantic generation stage, visual features are extracted by ViT (Dosovitskiy2020Animage), projected into a unified semantic space via Q-Former (Li2023BLIP-2), and embedded into prompts with structural tags. The full sequence is then processed by Qwen2.5-VL-Instruct (Bai2023Qwen-vl) to generate descriptions. A case example is shown in Figure 3 (see Section Case Study).
@@
where $\mathbf{E}_{\text{text}}$ is obtained from the [CLS] token of Chinese-RoBERTa-wwm-ext (Sun2021ChineseBERT); $\mathbf{E}_{\text{lvl}}$ encodes the heading path $\{h_{1},h_{2},h_{3},h_{4}\}$ via a GRU:
@@
Pharos-ESG is compared to three system groups: (1) Dedicated document parsers: MinerU (Huang2022LayoutLMv3), Marker (Tito2023Hierarchical), Docling (Kim2023Donut), and Textin (Appalaraju2021DocFormer), evaluated on reading order prediction, hierarchy alignment, and structural parsing. (2) General-purpose multimodal models: GPT-4o, Gemini 2.5 Pro, Doubao, DeepSeek-V3, DeepSeek-R1, and Qwen 3, evaluated on ToC Extraction and tasks from the first group. (3) Traditional and neural classifiers: SVM+TF-IDF (Joachims2005Text), XGBoost (Chen2016XGBoost), BERT-base (Devlin2019BERT), HAN (Huang2021Hierarchical), and HMCN (Sadat2022Hierarchical), evaluated on multi-level label prediction.
@@
  • Reading order prediction: Evaluated using Reading Order Kendall’s Tau (ROKT), a reading-order consistency metric derived from Kendall’s Tau (kendall1938new).

### v2

@@ Introduction
These challenges have driven financial research to adopt indirect proxies for ESG content, such as simple disclosure indicators (Ni and Zhang 2019), small-scale case studies (Florian Berg 2022), or third-party ratings (Dane M. Christensen 2022), all of which bypass the rich semantics of the reports themselves. Meanwhile, existing document parsers, designed for structurally regular formats like academic papers, legal contracts, or forms (Palm et al. 2019; Xu et al. 2020; Huang et al. 2022), struggle with the irregularity and implicit structure of ESG reports, limiting their applicability to this domain.
To address these challenges, we present Pharos-ESG, a unified framework for multimodal parsing, contextual grounding, and hierarchical labeling of ESG reports (Figure 2). Pharos-ESG integrates four core components: (a) reading-order modeling based on layout flow, (b) structure reconstruction guided by table-of-contents (ToC) anchors, (c) contextual transformation of visual elements into natural language, and (d) multi-level labeling across ESG topics, GRI indicators, and sentiment. The system outputs reading-order-aligned structures with linked text, tables, and images, enriched by hierarchical and semantic annotations. Experiments on expert-annotated benchmarks demonstrate that Pharos-ESG consistently outperforms both specialized parsers (e.g., MinerU (Huang et al. 2022), Marker (Tito et al. 2023), Docling (Kim et al. 2023)) and general-purpose multimodal models (e.g., GPT-4o, Gemini 2.5 Pro, Doubao) across multiple tasks.
@@
Recent advances in document intelligence have significantly improved performance on general-purpose document understanding tasks (Xu et al. 2020; Huang et al. 2022). However, ESG reports remain challenging due to chaotic reading order from irregular, slide-like layouts and implicit hierarchies caused by lengthy, weakly structured content.
@@
Dedicated document parsing systems such as Docling (Livathinos et al. 2025), MinerU (Wang et al. 2024), Marker (Datalab 2025), and TextIn (intsig-textin 2023), are effective at converting documents into structured outputs. They often incorporate state-of-the-art layout analysis models like EfficientViT (Liu et al. 2023), RT-DETR (Zhao et al. 2024), and LayoutLMv3 (Huang et al. 2022), achieving high accuracy on datasets with regular layouts such as PubLayNet (Zhong et al. 2019). However, their performance degrades significantly on structurally diverse datasets like DocLayNet (Pfitzmann et al. 2022), revealing limitations in generalization to visually irregular documents, including ESG reports. In parallel, vision-language models (VLMs) such as GPT (OpenAI), Gemini (Google), and DeepSeek (DeepSeek) provide a flexible, end-to-end solution by directly converting page images into text. However, these models often hallucinate and incur substantial computational costs when reconstructing implicit hierarchies in long, weakly structured ESG reports.
@@
Transformer-based models such as DocFormer (Appalaraju et al. 2021), LayoutLMv3 (Huang et al. 2022), StrucTexT (Appalaraju et al. 2021), UDoc (Kim et al. 2023), and SelfDoc (Li et al. 2021) employ cross-modal attention to associate figures with captions and link textual blocks with related tables or legends. While effective on structurally regular documents, they struggle with ESG reports, where semantically related elements are often dispersed across non-adjacent, even cross-page regions.
@@
Traditional approaches typically model structure at the page level, limiting global context and failing to capture cross-page semantic and hierarchical links (Huang et al. 2022; Xia et al. 2022). Most rely on explicit visual cues (e.g., numbering, indentation, font size) which often prove unreliable in complex or inconsistent layouts (Wang et al. 2023; Zhong et al. 2019; Xu et al. 2020). Recent models combine multimodal encoders with structure-aware decoders for joint hierarchy prediction, such as HRDoc (Marulanda et al. 2023) and DINO (Zhang et al. 2023). Yet, attention decay in lengthy documents (e.g., 20+ pages) still leads to recognition failures (Beltagy et al. 2020; Huang et al. 2022). These challenges are amplified in ESG reports, which feature exceptional length, weak structural cues, and diverse formatting across issuers.
@@
To our knowledge, four recent datasets support ESG-related research. Morio and Manning (2023) released 10,000+ corporate climate policy documents. ESG-FTSE (Wang and Casey 2024) contains 3,913 news articles on the UK’s top 10 firms. A3CG (Ong et al. 2025) compiles ESG reports from 1,679 Singapore-listed companies, and DynamicESG (Tseng et al. 2023) includes 2,220 Taiwan-based ESG news articles. However, these datasets offer limited coverage of full-length, multimodal ESG reports and lack the fine-grained annotations required for downstream financial tasks.
@@
Based on the extraction of multimodal elements from ESG reports via a structured preprocessing pipeline—including metadata extraction (Han et al. 2003), layout analysis (Huang et al. 2022), content parsing (Zhang et al. 2020), and noise reduction (JaidedAI 2024)—Pharos-ESG adopts and extends the Immediate Succession During Reading paradigm (Zhang et al. 2024) for reading order modeling, transitioning from global sequence ranking to a pairwise succession classification framework.
@@
where $E(w_{i})$ and $E(w_{j})$ are content embeddings: text and tables are encoded using LayoutLMv3 (Huang et al. 2022) For image blocks, the local image URL is used only as a placeholder for block identity and is not encoded nor used as a visual feature in the model. $\Delta y_{ij}$ and $\Delta x_{ij}$ denote center offsets; $\mathrm{IoU}{ij}$ and $\mathrm{Dist}{ij}$ are the bounding-box overlap and normalized distance; and $E(c_{i}),E(c_{j})$ are one-hot encodings of block types.
@@
To align the hierarchical structure parsed from ToC with the corresponding content in the document body, we propose ALIGN (Anchor-based Linguistic Indexing for Granular Navigation), a multi-stage alignment algorithm designed for visually heterogeneous layouts. ALIGN operates in three stages: (1) exact match, performing character-level matching to identify high-confidence anchors; (2) fuzzy/containment matching, using Levenshtein similarity (Mehlhorn and Näher 1990) and substring containment to expand coverage; and (3) context-aware insertion, resolving unmatched ToC headings by reasoning within anchor-defined windows in the document body. To this end, a large language model is prompted using a tailored CIP (Context-aware Insertion Prompt) to perform four structured inference steps: (i) summarizing paragraphs to generate semantic representations; (ii) assessing whether the region lacks an overview heading; (iii) identifying where the unmatched heading improves local structure; and (iv) selecting the optimal insertion point based on structural and semantic alignment.
@@
where $h_{i}$ is the hierarchical heading path, $\{x_{1},\ldots,x_{k}\}$ are the ordered elements, and $q_{i}$ is the task instruction. The prompt comprises (i) heading path embedding, linking the cluster to its section context, and (ii) element declaration templates, specifying type, order, and content. In the multimodal embedding and semantic generation stage, visual features are extracted by ViT (Dosovitskiy et al. 2020), projected into a unified semantic space via Q-Former (Li et al. 2023), and embedded into prompts with structural tags. The full sequence is then processed by Qwen2.5-VL-Instruct (Bai et al. 2023) to generate descriptions. A case example is shown in Figure 3 (see Section Case Study).
@@
where $\mathbf{E}_{\text{text}}$ is obtained from the [CLS] token of Chinese-RoBERTa-wwm-ext (Sun et al. 2021); $\mathbf{E}_{\text{lvl}}$ encodes the heading path $\{h_{1},h_{2},h_{3},h_{4}\}$ via a GRU:
@@
Pharos-ESG is compared to three system groups: (1) Dedicated document parsers: MinerU (Huang et al. 2022), Marker (Tito et al. 2023), Docling (Kim et al. 2023), and Textin (Appalaraju et al. 2021), evaluated on reading order prediction, hierarchy alignment, and structural parsing. (2) General-purpose multimodal models: GPT-4o, Gemini 2.5 Pro, Doubao, DeepSeek-V3, DeepSeek-R1, and Qwen 3, evaluated on ToC Extraction and tasks from the first group. (3) Traditional and neural classifiers: SVM+TF-IDF (Joachims 2005), XGBoost (Chen and Guestrin 2016), BERT-base (Devlin et al. 2019), HAN (Huang et al. 2021), and HMCN (Sadat and Caragea 2022), evaluated on multi-level label prediction.
@@
  • Reading order prediction: Evaluated using Reading Order Kendall’s Tau (ROKT), a reading-order consistency metric derived from Kendall’s Tau (Kendall 1938).